# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v4

@@ Results.
Figure 6 presents the nonlinear MCC results with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are consistently excluded from the representations (MCC $=0$). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in § D.2.
Figure 5: Random samples from MPI3D-Complex. Images are captured in a controlled setting.
@@
5.3 Causal3DIdent Dataset
@@
We further conduct experiments on the Causal3DIdent dataset [80, 66, 12, 76], a semi-synthetic dataset allowing explicitly defined causal relations among latent variables. Images are generated from 10 latent variables: 3 discrete (shape, x_pos, y_pos) and 7 continuous (color, s_pos, s_color, b_color, alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific, and the remaining as semantic variables following a causal graph among latent variables (Figure 15). See Figure 7 for example images from the dataset.
[…]
Training paradigm mirrors that of MPI3D-Complex. We evaluate using $R^{2}$ for continuous and MCC for discrete latent variables, averaged over three seeds. See § E.1 for full experimental details.
@@ Results.
Figure 8 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables shared across modalities, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably recovered by MMCL across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$.
Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure.
Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in § E.2.
Figure 7: Random samples from Causal3DIdent. Images are generated from the graph in Figure 15.
@@
Figure 8: Predicting semantic variables under misalignment using image features. $R^{2}$ is reported for continuous factors and MCC for discrete factors. Left: selection bias. Right: perturbation bias.
6 Related Work
Theoretical Multimodal (Multi-View) Contrastive Learning.
Vision-Language Models and Perspectives on Misalignment.
Identifiability in Latent Variable Models.
Invariant Representation Learning.
7 Conclusion and Discussion
@@
In this work, we present a formal analysis of cross-modal misalignment in multimodal data pairs within the MMCL framework, examining its impact on the learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, while discarding misaligned latent variables. When image-text pairs exhibit misalignment due to selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights highlight the need for multimodal learning frameworks that mitigate misalignment or leverage beneficial biases to enhance representation learning in real-world settings. More discussion of these implications is provided in App. G.
@@ Appendix A Notation and Terminology
Table 1 provides a summary of the notations and terminologies used throughout the paper.
@@
$\mathbb{I}_{\mathbf{s}}$ Index set of semantic variables: $[n_{s}]$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq. 1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq. 2)
@@
$\theta$ Selection bias, an integer realization in the range $[2^{n_{s}}-1]$
$\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ The set of all non-empty subsets of $\mathbb{I}_{\mathbf{s}}$
$\mathbb{I}_{\theta}$ A selected semantic subset indexed by $\theta$, $\mathbb{I}_{\theta}\in\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$
@@
$\rho$ Perturbation bias, an integer realization in the range $[2^{|\mathbb{I}_{\theta}|}-1]$
$\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$ The set of all proper subsets of $\mathbb{I}_{\theta}$
$\mathbb{I}_{\rho}$ A subset of $\mathbb{I}_{\theta}$ subject to perturbation indexed by $\rho$, $\mathbb{I}_{\rho}\in\mathcal{P}_{\text{proper}}(\mathbb{I}_{\theta})$
@@
$p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}$ Perturbation conditional distribution, reflecting misalignment (Eq. 5)
@@
Appendix B Proofs
B.1 Lemmas
@@
Lemma B.1 (Global Minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$).
In the context of the proposed LVM as outlined in § 3.1, and Asms. 3.1 and 3.2, the global minimum of
@@
  • $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ denote the true underlying generative mappings for images and paired text, respectively, as described in § 3.
@@
Proof of Lem. B.1.
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations 7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq. 6, we have
@@
By the invertibility of the generative processes $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ (see § 3), we may change variables to express the expectation over the latent variables:
@@
We now show that these unbiased semantic components are identical across modalities almost everywhere (a.e.). By Asm. 3.2, for any image-text pair the text is generated via a random perturbation process that modifies only a subset $A\subseteq\mathbb{I}_{\rho}$ of the activated semantic variables. Specifically, recall Eq. 5, the perturbation density is defined as
@@
Lemma B.2 (Uniformizing Mapping Preserves All Information).
Let $\mathbf{h}:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=\mathbf{h}(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $\mathbf{h}$ is a global diffeomorphism; in other words, $\mathbf{h}$ is bijective and depends on every component of $\mathbf{u}$ [^7] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
Proof of Lem. B.2.
@@
B.2 Proof of Theorem 4.1
We now proceed to prove Thm. 4.1. To begin, we restate the theorem for clarity:
@@
    First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 00. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
Since all terms in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ are non-negative, and its global minimum is 00 by Lem. B.1, each term in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. Thus, minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ leads to:
@@
From Eq. 9, it follows that
@@
which ensures alignment of representations a.s. for any pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. The substitution of expectations for the image modality in Eq. 9 is valid because $\mathbf{x}$ follows the pushforward distribution of $\mathbf{z}_{x}$ under the deterministic diffeomorphism $\mathbf{g}_{x}$ (Eq. 3); a similar argument applies to the text (Eq. 4).
Equations 10 and 11 imply that $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ onto uniform distributions over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$), since their differential entropy equals to zero.
@@
Since $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$ are the latent variables generating images $\mathbf{x}$ and paired text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, respectively, the data-generating process in § 3 implies the following independence properties:
@@
Combining Equations 12 and 13, we have:
@@
Since $\mathbf{m}_{x}$ is independent of $\mathbf{s}$, perturbations in $\mathbf{m}_{x}$ does not alter the semantic variables, and $p_{\mathbf{m}_{x}}>0$ a.e. over $\mathcal{M}_{x}$ by Asm. 3.1. Thus, substituting
@@
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$ with $\boldsymbol{\varsigma}$ a small enough perturbation, then by Eq. 14 in statement (c2), $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq. 12 then forces $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $\mathbf{h}_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
@@
Using the invariance condition in Eq. 12, together with the independence of modality-specific non-semantic variables in Equations 17 and 18, we have the following updated invariance condition:
@@
However, combining Eq. 19, we have
@@
Equations 21 and 22 thus contradict each other. Hence, such a function $\mathbf{h}^{c}_{x}$ cannot exist. Consequently, $\mathbf{h}_{x}$ must be independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
@@
Clarification B.1 (Causal Interpretations).
(i) Justification for the existence of distinct points in Eq. 20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm. 3.1. From a latent SCM perspective [51, 65], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
First, we refine the invariance condition by excluding omitted semantic variables as established above. Combining Equations 12, 18 and 23, we obtain:

### v5

@@ Results.
Figure˜6 presents the nonlinear MCC results with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are consistently excluded from the representations (MCC =0=0). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in §˜E.2.
@@
5.3 Causal3DIdent: Semi-Synthetic Dataset with Structured Causal Latent Variables
@@
Figure 7: Causal3DIdent samples.
We further conduct experiments on the Causal3DIdent dataset [83, 68, 13, 78], a semi-synthetic dataset allowing explicitly defined causal relations among latent variables. Images are generated from 10 latent variables: 3 discrete (shape, x_pos, y_pos) and 7 continuous (color, s_pos, s_color, b_color, alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific, and the remaining as semantic variables following a causal graph among latent variables (Figure˜17). See Figure˜7 for example images from the dataset. For text, we discretize color-based attributes (color, s_color, b_color), retain continuous s_pos and simulate partial loss in spotlight information when generating text. Text descriptions are rendered using five templates, based on latent semantic variables under selection and perturbation bias. Specifically, selection bias settings ($\mathbb{I}_{\theta}$) vary from {shape} to the set of all seven semantic indices, with $\mathbb{I}_{\rho}=\emptyset$; perturbation bias settings progressively perturb subsets $\mathbb{I}_{\rho}$ from $\emptyset$ to {x_pos, y_pos, s_pos, color, s_color, b_color} in reverse order, with full semantic selection.
Training paradigm mirrors that of MPI3D-Complex. We evaluate using $R^{2}$ for continuous and MCC for discrete latent variables, averaged over three seeds. See §˜F.1 for full experimental details.
@@ Results.
Figure˜8 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables shared across modalities, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably recovered by MMCL across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$. Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure. Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in §˜F.2.
@@
Figure 8: Predicting semantic variables under misalignment using image features. $R^{2}$ is reported for continuous factors and MCC for discrete factors. The predictions pattern align with our theoretical results, though extra dimensions may be predictable due to statistical correlations.
5.4 Case Study: Zero-Shot Evaluation of OpenCLIP Model Representations
We further validate our theoretical findings through a comprehensive zero-shot evaluation case study on OpenCLIP models [24], a foundation MMCL model pretrained on the LAION-400M dataset [61]. To achieve this, we develop a structured taxonomy comprising 146 visual concepts, organized into 15 distinct concept groups with corresponding abbreviations for clarity in plots and analyses: Animal, Clothing, Color, Food, Object, Role (i.e., human roles), Scene, Vehicle, Weather, Texture, POV (i.e., point of view of the photo), Emot. (i.e., emotion or mental state), Postproc. (i.e., post-processing), Trait (i.e., trait judgment by appearance), and Stere. (i.e., stereotypes). For each concept, we collect a dataset of up to 200 images from Flickr [^7] https://www.flickr.com/ by prompting with concept names, available under CC licenses. See the concept taxonomy and data collection details in §˜G.1.
Concept Coverage in LAION-400M Captions.
We begin by analyzing the concept coverage of LAION-400M captions, measuring the average coverage rate, defined as the percentage of captions that explicitly mention each concept or its synonyms, averaged by concept group. As shown in Figure˜9, concept groups display substantial variation in average coverage: common concepts exhibit relatively high coverage (e.g., Object: 1.63%, Color: 2.16%), while some concepts are rarely mentioned (e.g., Stere.: 0.0003%, Postproc.: 0.0118%). Importantly, due to the large scale of the dataset, no concept group exhibited 0% coverage, indicating that complete omission by selection bias is unlikely. Thus, the coverage rate serves as an approximate indicator of selection bias across concept groups.
Zero-Shot Evaluation of OpenCLIP Models.
Figure˜10 shows the zero-shot F1 performance of OpenCLIP across concept groups. Consistent with our theoretical findings, OpenCLIP models, regardless of scale, perform well on frequently captioned groups (e.g., Animal, Object), but their performance declines sharply for under-captioned groups (e.g., Trait, Emot.). Concept-level confusion matrices further reveal systematic misclassifications, particularly within low-coverage groups (see §˜G.3). Notably, omitting captions for sensitive concepts such as Stere. may be intentional to avoid biased knowledge. In contrast, under-captioned yet valuable concepts like Texture and Emot. should be learned, suggesting a need for more targeted captioning strategies.
Figure 9: Concept coverage by group in LAION-400M captions. See §˜G.2 for intra-group details.
Figure 10: Zero-shot F1 score of OpenCLIP model trained on LAION-400M in each concept group.
6 Conclusion
@@
In this work, we present a formal analysis of cross-modal misalignment in multimodal data pairs within the MMCL framework, examining its impact on the learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, while discarding misaligned latent variables. When image-text pairs exhibit misalignment due to selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights highlight the need for multimodal learning frameworks that mitigate misalignment or leverage beneficial biases to enhance representation learning in real-world settings. More discussion of these implications is provided in App.˜I.
@@ Appendix A Notation and Terminology
Table˜1 provides a summary of the notations and terminologies used throughout the paper.
@@
$\mathbb{I}_{\mathbf{s}}$ Index set of semantic variables: $n_{s}]$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq.˜1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq.˜2)
@@
$\theta$ Selection bias, an integer realization in the range $2^{n_{s}}-1]$
$\mathcal{P}^{+}(\mathbb{I}_{\mathbf{s}})$ The set of all non-empty subsets of $\mathbb{I}_{\mathbf{s}}$
$\mathbb{I}_{\theta}$ A selected semantic subset indexed by $\theta$, $\mathbb{I}_{\theta}\in\mathcal{P}^{+}(\mathbb{I}_{\mathbf{s}})$
@@
$\rho$ Perturbation bias, an integer realization in the range $2^{|\mathbb{I}_{\theta}|}-1]$
$\mathcal{P}_{\text{prop}}(\mathbb{I}_{\theta})$ The set of all proper subsets of $\mathbb{I}_{\theta}$
$\mathbb{I}_{\rho}$ A subset of $\mathbb{I}_{\theta}$ subject to perturbation indexed by $\rho$, $\mathbb{I}_{\rho}\in\mathcal{P}_{\text{prop}}(\mathbb{I}_{\theta})$
@@
$p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho}$ Perturbation conditional distribution, reflecting misalignment (Eq.˜5)
@@
Appendix B Related Work
Theoretical Multimodal (Multi-View) Contrastive Learning.
Vision-Language Models and Perspectives on Misalignment.
Identifiability in Latent Variable Models.
Invariant Representation Learning.
Appendix C Proofs
C.1 Lemmas
@@
Lemma C.1 (Global Minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$).
In the context of the proposed LVM as outlined in §˜3.1, and Asms.˜3.1 and 3.2, the global minimum of
@@
  • $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ denote the true underlying generative mappings for images and paired text, respectively, as described in §˜3.
@@
Proof of Lem.˜C.1.
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations˜7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq.˜6, we have
@@
By the invertibility of the generative processes $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ (see §˜3), we may change variables to express the expectation over the latent variables:
@@
We now show that these unbiased semantic components are identical across modalities almost everywhere (a.e.). By Asm.˜3.2, for any image-text pair the text is generated via a random perturbation process that modifies only a subset $A\subseteq\mathbb{I}_{\rho}$ of the activated semantic variables. Specifically, recall Eq.˜5, the perturbation density is defined as
@@
Lemma C.2 (Uniformizing Mapping Preserves All Information).
Let $\mathbf{h}:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=\mathbf{h}(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $\mathbf{h}$ is a global diffeomorphism; in other words, $\mathbf{h}$ is bijective and depends on every component of $\mathbf{u}$. [^8] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
Proof of Lem.˜C.2.
@@
C.2 Proof of Theorem 4.1
We now proceed to prove Thm.˜4.1. To begin, we restate the theorem for clarity:
@@
    First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 0. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
Since all terms in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ are non-negative, and its global minimum is 0 by Lem.˜C.1, each term in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. Thus, minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ leads to:
@@
From Eq.˜9, it follows that
@@
which ensures alignment of representations a.s. for any pair $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$. The substitution of expectations for the image modality in Eq.˜9 is valid because $\mathbf{x}$ follows the pushforward distribution of $\mathbf{z}_{x}$ under the deterministic diffeomorphism $\mathbf{g}_{x}$ (Eq.˜3); a similar argument applies to the text (Eq.˜4).
Equations˜10 and 11 imply that $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ onto uniform distributions over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$), since their differential entropy equals to zero.
@@
Since $\mathbf{z}_{x}=(\mathbf{s},\mathbf{m}_{x})$ and $\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}=(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$ are the latent variables generating images $\mathbf{x}$ and paired text $\mathbf{t}^{\scriptscriptstyle(\theta)}$, respectively, the data-generating process in §˜3 implies the following independence properties:
@@
Combining Equations˜12 and 13, we have:
@@
Since $\mathbf{m}_{x}$ is independent of $\mathbf{s}$, perturbations in $\mathbf{m}_{x}$ does not alter the semantic variables, and $p_{\mathbf{m}_{x}}>0$ a.e. over $\mathcal{M}_{x}$ by Asm.˜3.1. Thus, substituting
@@
A symmetric argument applies to $\mathbf{m}_{t}$. If $\mathbf{m}_{t}\mapsto\mathbf{m}_{t}+\boldsymbol{\varsigma}$ with $\boldsymbol{\varsigma}$ a small enough perturbation, then by Eq.˜14 in statement (c2), $\mathbf{h}_{x}(\mathbf{s},\mathbf{m}_{x},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ remains unchanged. The invariance condition in Eq.˜12 then forces $\mathbf{h}_{t}(\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}},\mathbf{m}_{t}+\boldsymbol{\varsigma})$ to remain constant w.r.t. $\boldsymbol{\varsigma}$, showing that $\mathbf{h}_{t}$ is independent of $\mathbf{m}_{t}$. Therefore, the learned representations satisfy:
@@
Using the invariance condition in Eq.˜12, together with the independence of modality-specific non-semantic variables in Equations˜17 and 18, we have the following updated invariance condition:
@@
However, combining Eq.˜19, we have
@@
Equations˜21 and 22 thus contradict each other. Hence, such a function $\mathbf{h}^{c}_{x}$ cannot exist. Consequently, $\mathbf{h}_{x}$ must be independent of $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$. Formally,
@@
Clarification C.1 (Causal Interpretations).
(i) Justification for the existence of distinct points in Eq.˜20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm.˜3.1. From a latent SCM perspective [53, 67], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
First, we refine the invariance condition by excluding omitted semantic variables as established above. Combining Equations˜12, 18 and 23, we obtain: